# Can We Predict Before Executing Machine Learning Agents?

### v1

@@ Abstract
Autonomous machine learning agents have revolutionized scientific discovery, yet they remain constrained by a Generate-Execute-Feedback paradigm. Previous approaches suffer from a severe Execution Bottleneck, as hypothesis evaluation relies strictly on expensive physical execution. To bypass these physical constraints, we internalize execution priors to substitute costly runtime checks with instantaneous predictive reasoning, drawing inspiration from World Models. In this work, we formalize the task of Data-centric Solution Preference and construct a comprehensive corpus of 18,438 pairwise comparisons. We demonstrate that LLMs exhibit significant predictive capabilities when primed with a Verified Data Analysis Report, achieving 61.5% accuracy and robust confidence calibration. Finally, we instantiate this framework in ForeAgent, an agent that employs a Predict-then-Verify loop, achieving a 6x acceleration in convergence while surpassing execution-based baselines by +6%. Our code and dataset will be publicly available soon at https://github.com/zjunlp/predict-before-execute.
@@ 3.2 Source and Scope
To instantiate the task inputs defined above, we construct a large-scale corpus derived from the real-world execution trajectories of two ML agents, AIDE Jiang et al. (2025) and AutoMind Ou et al. (2025), operating on MLE-bench Chan et al. (2025) platform (Figure 2(b)). Powered by DeepSeek-V3.1 DeepSeek-AI (2025b) and o3-mini OpenAI (2025a), these agents generate 1,329 valid solutions across 26 diverse tasks (Table 1). Unlike synthetic snippets, these candidates represent complete ML workflows ranging from preprocessing to training. Therefore, identifying the superior solution requires evaluating how well an algorithm fits the specific data characteristics, rather than merely checking for code syntax.
@@ 3.3 Dataset Curation and Instantiation
To ensure rigorous evaluation, we implement an Expert-in-the-Loop pipeline to prune raw trajectories into 895 high-quality instances. This process involves deduplication, automated taxonomy tagging, and expert sampling to cap dominant methods and ensure algorithmic diversity. Subsequently, we instantiate the dataset by exhaustively generating pairwise combinations from this curated corpus. We apply strict filtering to discard ambiguous pairs and balance the ground-truth winner’s position to mitigate position bias Shi et al. (2024). This yields a final dataset of 18,438 comparisons (Figure 2(c)), utilizing micro-averaged accuracy as the primary metric.
@@ 3.4 Input Augmentation: The Verified Data Analysis Report
To address LLMs’ numerical limitations Davies et al. (2025); Li et al. (2025b) and context constraints preventing direct data ingestion, we augment inputs with a Verified Data Analysis Report that transforms raw statistics into semantic narratives Rytting and Wingate (2021); Zhang et al. (2025a). To guarantee factual grounding, we implement a strict “Code-Execution-Verbalization” protocol (Figure 2(d)). GPT-5.1 OpenAI (2025b) first performs Code Profiling on raw data files (strictly masking labels and outcomes) to generate analysis scripts; these scripts undergo Execution & Verification to yield artifact-free logs; finally, GPT-5.1 performs Verbalization to translate these logs into the Verbal Data Report (Data Observation $\to$ Modeling Implication), ensuring reliable semantic grounding for the task (see case in Appendix Figure 10).
@@ 5 Analysis & Insights
Figure 3: Comprehensive Analysis of World Model Mechanisms and Capabilities. (a) Impact of Data Representation: Predictive success stems from semantic data understanding rather than complexity heuristics. (b) Domain Sensitivity: The superiority of verbal reports remains consistent across domains. (c) Scaling Laws: Accuracy decouples from pure parameter scaling. (d) Inference Dynamics: Active reasoning outperforms direct answering with robust stability across temperatures. (e) Calibration Analysis: Self-reported confidence strictly correlates with accuracy. (f) Complexity Discrimination: Accuracy scales with the complexity gap.
@@ 5.2 RQ2: Capabilities, Boundaries, and Algorithmic Bias
Figure 3(d) identifies reasoning as the primary engine, with the Thinking Mode (CoT) DeepSeek-AI (2025a); OpenAI (2024) ($61.3\%$) outperforming Direct Answering ($55.9\%$). This performance remains robust across temperatures ($T\in[0,1.5]$), implying an invariant logical core despite diversified traces. However, this capability is constrained by the problem landscape; Table 2 reveals sharp performance stratifications across the Task-Solution matrix. On the Task Dimension, models demonstrate a preference for NLP ($66.9\%$) and Easy ($63.9\%$) paradigms. Simultaneously, the Solution Dimension reveals a “Complexity Tax” ($59.6\%$ on complex code) and a granularity bottleneck, where the model is more effective at distinguishing broad Cross-Algo contrasts (comparing solutions with different algorithms, $62.8\%$). Thus, while reasoning is indispensable, it faces limits when navigating intricate code logic or subtle intra-class nuances.
@@
Tracing the Complexity Gap in Figure 3(f) shows accuracy scales with distinction; crucially, the model’s superiority over heuristics in low-gap scenarios proves it detects valid semantic signals rather than simple metrics. Furthermore, Figure 3(e) demonstrates excellent Calibration, where confidence correlates strictly with accuracy. This reliability underpins the Section 6 gating mechanism, ensuring agents act with certainty.
@@ 6.2 Method: The Predict-then-Verify Loop
We propose ForeAgent, which re-engineers the Improvement stage into a conservative Predict-then-Verify loop (Figure 2(e)) to bridge the Implementation Gap Zhu et al. (2025a). The workflow proceeds through three key phases: (1) High-Volume Generation, where $m=10$ candidates are proposed in parallel to expand search width without execution costs; (2) Confidence-Gated Pairwise Selection, which utilizes a confidence gate ($c=0.7$) to ensure high-certainty selection; and (3) Verification Execution, where the Top-$k$ ($k=1$) candidate is physically verified to anchor the solution trajectory in execution feedback.
@@ 6.4 Results
By substituting costly execution with rapid inference, ForeAgent achieves an average $\bm{6\times}$ speedup (Figure 4(b)), enabling it to explore $\bm{3.2\times}$ more nodes within just 1/6 of the time budget (Figure 4(c)). This expanded search capability directly translates into performance, driving a $\bm{+6\%}$ improvement in Beat Ratio (Figure 4(a)) and demonstrating robust generalization on unseen tasks. Although we currently focus on inference, this paradigm naturally extends to training contexts like Reward Model, a promising direction we reserve for future work.
@@ 8 Conclusion
Although our corpus encompasses 18,438 pairs across 26 tasks, the distribution remains inherently skewed. Mainstream paradigms like Classification and Regression dominate the dataset, whereas niche scientific tasks (e.g., Audio Classification, Tabular Grading) are represented by significantly smaller sample sizes. Consequently, while the model demonstrates strong general capabilities, its reliability in extremely low-resource or highly specialized scientific domains may vary, and the current evaluation may not fully reflect the challenges of these long-tail scenarios.
@@ Appendix A Extended Related Work
LLMs with strong reasoning capabilities Qiao et al. (2023) are increasingly serving as core controllers for autonomous agents in scientific discovery Gu et al. (2024b); Chen et al. (2025c), extending to specialized machine research domains Toledo et al. (2025); Zhang et al. (2025d). Beyond the digital realm, agents are transforming laboratory research Liu et al. (2025a); Li et al. (2025d); Huang et al. (2025); Schmidgall et al. (2025) and complex data analytics Sun et al. (2025b); Zhang et al. (2025c). Prominent systems now autonomously propose hypotheses Chai et al. (2025); Yu et al. (2025b); Team et al. (2025); Novikov et al. (2025) and conduct closed-loop experiments Gottweis et al. (2025); Lange et al. (2025); Yuan et al. (2025); Yu et al. (2025a), highlighting the trend of “AI Scientists” operating in open-ended exploration loops.
Narrowing down to the machine learning domain, the ecosystem is highly diversified. One stream of research focuses on managing the end-to-end workflow, ranging from autonomous frameworks Nam et al. (2025); Yang et al. (2025); Qiao et al. (2025); Ze-xi et al. (2025) to engineering pipelines Fang et al. (2025); Chi et al. (2024). Another stream, driven by benchmarks like MLE-bench Chan et al. (2025); Huang et al. (2024) and broader evaluation suites Zhang et al. (2025e); Jing et al. (2024); Zhang et al. (2024b); Nathani et al. (2025), focuses on competitive problem-solving through knowledge-guided reasoning Luo et al. (2025); Ou et al. (2025) and evolutionary optimization Du et al. (2025); Guo et al. (2024); Li et al. (2025a); Liu et al. (2025b).
Additionally, general-purpose platforms and optimization frameworks offer the foundational tooling and multi-agent architectures required for scalable research Wang et al. (2025b); Jiang et al. (2025); Hong et al. (2024); Qiang et al. (2025); Wang et al. (2025a). However, to mitigate the significant computational overhead of the generation-execution-feedback loop inherent in these systems, recent approaches explore utilizing internal priors to estimate feasibility and prune redundant steps, thereby accelerating optimization Kulibaba et al. (2025); Trirat et al. (2025); Zhang et al. (2024a); Astorga et al. (2025).
@@
  • AIDE: AIDE (Jiang et al., 2025) is an LLM-based agent that frames machine learning engineering as a code optimization problem. It structures the trial-and-error process as a tree search in the solution space, systematically reusing and refining promising code candidates. This method effectively trades computational resources for enhanced performance. Specifically, AIDE first generates initial code $C_{0}$ based on instruction $I$. The code is executed by training on dataset $D$ to obtain results. Subsequently, AIDE iteratively derives new code $C_{1},C_{2},\dots,C_{t}$ based on the feedback.
@@
Recent research enables models to internalize the execution process, predicting test outcomes Hora (2024); team et al. (2025); Wei et al. (2025b) or assessing logic consistency directly Li et al. (2025c); Chang-shu et al. (2024). This capability is rigorously evaluated on reasoning-centric benchmarks Wei et al. (2025a); Gu et al. (2024a); Jain et al. (2024). Unlike traditional benchmarks that may allow for rote memorization, these tasks require models to transcend statistical pattern matching and develop a deep semantic understanding of algorithmic states and control flows Chen et al. (2025d); Sun et al. (2025a); Akhauri et al. (2025); Chen et al. (2025a), serving as the foundational capability for our proposed framework. Aligning with OpenAI’s Level 4 “Innovators” Metz (2024); Zhang et al. (2025b), this empowers agents to drive innovation by leveraging internal world models to proactively prune vast hypothesis spaces, shifting the paradigm from ensuring syntactic correctness to optimizing for semantic success. This transition resonates with the broader Data-Centric AI movement Zha et al. (2023); Cabrera et al. (2025), moving beyond model architecture to focus on the quality of evaluative signals. Specifically, our framework incorporates rationale-based preference optimization Just et al. (2024) and rigorous dataset construction criteria Shen et al. (2024) to ensure that the “implicit world model” is grounded in data-specific realities rather than abstract heuristics Tschalzev et al. (2024).
@@ Appendix B Corpus Details
To support the reproducibility of our analysis and provide a comprehensive view of the solution space, we provide detailed metadata for the prediction corpus.

### v2

@@ Abstract
Autonomous machine learning agents have revolutionized scientific discovery, yet they remain constrained by a Generate-Execute-Feedback paradigm. Previous approaches suffer from a severe Execution Bottleneck, as hypothesis evaluation relies strictly on expensive physical execution. To bypass these physical constraints, we internalize execution priors to substitute costly runtime checks with instantaneous predictive reasoning, drawing inspiration from World Models. In this work, we formalize the task of Data-centric Solution Preference and construct a comprehensive corpus of 18,438 pairwise comparisons. We demonstrate that LLMs exhibit significant predictive capabilities when primed with a Verified Data Analysis Report, achieving 61.5% accuracy and robust confidence calibration. Finally, we instantiate this framework in ForeAgent, an agent that employs a Predict-then-Verify loop, achieving a 6x acceleration in convergence while surpassing execution-based baselines by +6%. Our code and dataset are publicly available at https://github.com/zjunlp/predict-before-execute.
@@ 3.2 Source and Scope
To instantiate the task inputs defined above, we construct a large-scale corpus derived from the real-world execution trajectories of two ML agents, AIDE Jiang et al. (2025) and AutoMind Ou et al. (2025), operating on MLE-bench Chan et al. (2025) platform (Figure 2(b)). Powered by DeepSeek-V3.1 DeepSeek-AI (2025b) and o3-mini OpenAI (2025a), these agents generate 1,329 valid solutions across 26 diverse tasks (Table 1). Unlike synthetic snippets, these candidates represent complete ML workflows that are entirely generated by agents and absent from any pre-training corpora, heavily incorporating logically incomplete but executable intermediate states to model noisy real-world exploration (see Appendix B.4). Therefore, identifying the superior solution requires evaluating how well an algorithm fits the specific data characteristics, rather than merely checking for code syntax.
@@ 3.3 Dataset Curation and Instantiation
For rigorous evaluation, we implement an Expert-in-the-Loop pipeline to prune raw trajectories into 895 high-quality instances. This process involves deduplication, taxonomy tagging, and expert sampling to cap dominant methods and ensure algorithmic diversity. Next, we instantiate the dataset by exhaustively generating pairwise combinations from this curated corpus. We apply strict filtering to discard ambiguous pairs and balance the ground-truth winner’s position to mitigate position bias Shi et al. (2024). This yields a final dataset of 18,438 comparisons (Figure 2(c)), utilizing micro-averaged accuracy as the primary metric.
@@ 3.4 Input Augmentation: The Verified Data Analysis Report
To address LLMs’ numerical limitations Davies et al. (2025); Li et al. (2025b) and context constraints preventing direct data ingestion, we augment inputs with a Verified Data Analysis Report that transforms raw statistics into semantic narratives Rytting and Wingate (2021); Zhang et al. (2025a). To guarantee factual grounding and prevent hallucination, we implement a strict protocol (Figure 2(d)) consisting of three concrete steps: (1) Code Generation: GPT-5.1 OpenAI (2025b) generates a Python script to profile raw data with labels masked. For example, it writes “print(df[’target’].value_counts())” to inspect the target distribution. (2) Execution and Verification: The script runs in a sandbox to produce standard output. A human expert performs a strict pass or fail validity check to ensure the log is free of runtime errors. For instance, the execution log returns a raw fact like “Target Distribution: 0: 0.915, 1: 0.085”. (3) Verbalization: GPT-5.1 reads this execution log and translates it into a semantic insight. Following the previous example, it produces the final report: “Data Imbalance Warning: Severe class imbalance (Pos: 8.5%). Implication: Accuracy is not a suitable metric; consider using F1-score.” This process ensures reliable semantic grounding for the task (see case in Appendix Figure 12).
@@ 5 Analysis & Insights
Figure 3: Comprehensive Analysis of World Model Mechanisms and Capabilities. (a) Impact of Data Representation: Predictive success stems from semantic data understanding rather than complexity heuristics. (b) Domain Sensitivity: The superiority of verbal reports remains consistent across domains. (c) Scaling Laws: Accuracy decouples from pure parameter scaling. (d) Inference Dynamics: Active reasoning outperforms direct answering with robust stability across temperatures. (e) Calibration Analysis: Self-reported confidence correlates with accuracy. (f) Complexity Discrimination: Accuracy scales with the complexity gap.
@@ 5.2 RQ2: Capabilities, Boundaries, and Algorithmic Bias
Figure 3(d) identifies reasoning as the primary engine, with the Thinking Mode (CoT) DeepSeek-AI (2025a); OpenAI (2024) ($61.3\%$) outperforming Direct Answering ($55.9\%$). This performance remains robust across temperatures ($T\in[0,1.5]$) and trajectory variance (see Appendix C.2), implying an invariant logical core that relies on genuine reasoning rather than exploiting easy artifacts from the same agent run. However, this capability is constrained by the problem landscape; Table 2 reveals sharp performance stratifications across the Task-Solution matrix. On the Task Dimension, models demonstrate a preference for NLP ($66.9\%$) and Easy ($63.9\%$) paradigms. Simultaneously, analysis on the Solution Dimension reveals a comparison preference for Traditional ML within the Algo Era ($64.5\%$), a “Complexity Tax” ($59.6\%$ on complex code), and a Granularity bottleneck, where the model is more effective at distinguishing broad Cross-Algo contrasts (comparing solutions with different algorithms, $62.8\%$). Thus, while reasoning is indispensable, it faces limits when navigating intricate code logic or subtle intra class nuances.
@@
Tracing the Complexity Gap in Figure 3(f) shows accuracy scales with distinction; crucially, the model’s superiority over heuristics in low-gap scenarios proves it detects valid semantic signals rather than simple metrics. Furthermore, Figure 3(e) demonstrates Calibration, where confidence correlates strictly with accuracy. This reliability underpins the Section 6 gating mechanism, ensuring agents act with certainty.
@@ 6.2 Method: The Predict-then-Verify Loop
We propose ForeAgent, which re-engineers the Improvement stage into a conservative Predict-then-Verify loop (Figure 2(e)) to bridge the Implementation Gap Zhu et al. (2025a). The workflow proceeds through three key phases: (1) High-Volume Generation, where $m=10$ candidates are proposed in parallel to expand search width without execution costs; (2) Confidence-Gated Pairwise Selection, which utilizes a confidence gate ($c=0.7$) to ensure high-certainty selection; and (3) Verification Execution, where the Top-$k$ ($k=1$) candidate is physically verified to anchor the solution trajectory in execution feedback (see ablation study in Appendix C.6).
@@ 6.4 Results
By substituting costly execution with rapid inference, ForeAgent achieves an average $\bm{6\times}$ speedup (Figure 4(b)), enabling it to explore $\bm{3.2\times}$ more nodes within just 1/6 of the time budget (Figure 4(c)). This expanded search capability directly translates into performance, driving a $\bm{+6\%}$ improvement in Beat Ratio (Figure 4(a)) and demonstrating robust generalization on unseen tasks. Furthermore, the World Model acts as a semantic safeguard during intermediate development, significantly boosting the Test Improve Rate by 23% (see Appendix C.5). Although we currently focus on inference, this paradigm naturally extends to training contexts like Reward Model, a promising direction we reserve for future work.
@@ 8 Conclusion
Although our corpus encompasses 18,438 pairs across 26 tasks, the distribution remains inherently skewed. Mainstream paradigms like Classification and Regression dominate the dataset, whereas niche scientific tasks (e.g., Audio Classification, Tabular Grading) are represented by significantly smaller sample sizes. Consequently, while the model demonstrates strong general capabilities, its reliability in extremely low-resource or highly specialized scientific domains may vary, and the current evaluation may not fully reflect the challenges of these long-tail scenarios. Furthermore, the Verified Data Report currently relies on metadata for unstructured domains like CV and NLP, leaving the integration of multimodal data analysis agents for deeper semantic profiling to future work.
@@
Appendix Table of Contents
  • Appendix A: Extended Related Work
  • Appendix B: Corpus Details
    • B.1 Task Metadata and Scale
    • B.2 Algorithm and Architecture Distribution
    • B.3 Agent Evaluation Benchmark
    • B.4 Trajectory Sampling and Intermediate States
  • Appendix C: Detailed Experiment Result
    • C.1 Fine-grained Performance on Prediction Corpus
    • C.2 Analysis of Pair Source and Trajectory Variance
    • C.3 Detailed Performance Metrics of ForeAgent on AI4Science Benchmarks
    • C.4 Search Efficiency Analysis of ForeAgent
    • C.5 Decision Fidelity and Reliability in Local Iterations
    • C.6 Ablation Study on Top $k$ Selection
    • C.7 Licensing and Artifact Usage
    • C.8 Computational Infrastructure and Budget
    • C.9 Software Dependencies and Metric Implementation
  • Appendix D: Detailed Qualitative Analysis
    • D.1 Case I: Overcoming Complexity Bias
    • D.2 Case II: Domain Fit over Architectural Sophistication
    • D.3 Case III: Sample of the Verbal Data Report
    • D.4 Case IV: Sample of the Task Instruction ($I$)
  • Appendix E: Prompt Templates
@@ Appendix A Extended Related Work
LLMs with strong reasoning capabilities Qiao et al. (2023) are increasingly serving as core controllers for autonomous agents in scientific discovery Gu et al. (2024b); Chen et al. (2025c), extending to specialized machine research domains Toledo et al. (2025); Zhang et al. (2025d). Beyond the digital realm, agents are transforming laboratory research Liu et al. (2025a); Li et al. (2025d); Huang et al. (2025); Schmidgall et al. (2025) and complex data analytics Sun et al. (2025b); Zhang et al. (2025c). Prominent systems now autonomously propose hypotheses Chai et al. (2025); Yu et al. (2025b); Team et al. (2025); Novikov et al. (2025) and conduct closed-loop experiments Gottweis et al. (2025); Lange et al. (2025); Yuan et al. (2025); Yu et al. (2025a); Zhang et al. (2024a), highlighting the trend of “AI Scientists” operating in open-ended exploration loops.
Narrowing down to the machine learning domain, the ecosystem is highly diversified. One stream of research focuses on managing the end-to-end workflow, ranging from autonomous frameworks Nam et al. (2025); Yang et al. (2025); Qiao et al. (2025); Zexi et al. (2025) to engineering pipelines Fang et al. (2025); Chi et al. (2024); Zeng et al. (2025). Another stream, driven by benchmarks like MLE-bench Chan et al. (2025); Huang et al. (2024) and broader evaluation suites Zhang et al. (2025e); Jing et al. (2024); Zhang et al. (2024c); Nathani et al. (2025), focuses on competitive problem-solving through knowledge-guided reasoning Luo et al. (2025); Ou et al. (2025) and evolutionary optimization Du et al. (2025); Guo et al. (2024); Li et al. (2025a); Liu et al. (2025b).
Additionally, general-purpose platforms and optimization frameworks offer the foundational tooling and multi-agent architectures required for scalable research Wang et al. (2025b); Jiang et al. (2025); Hong et al. (2024); Qiang et al. (2025); Wang et al. (2025a). However, to mitigate the significant computational overhead of the generation-execution-feedback loop inherent in these systems, recent approaches explore utilizing internal priors to estimate feasibility and prune redundant steps, thereby accelerating optimization Kulibaba et al. (2025); Trirat et al. (2025); Zhang et al. (2024b); Astorga et al. (2025).
@@
  • AIDE: AIDE (Jiang et al., 2025) is an LLM-based agent that frames machine learning engineering as a code optimization problem. It structures the trial-and-error process as a tree search in the solution space, reusing and refining promising code candidates. This method effectively trades computational resources for enhanced performance. Specifically, AIDE first generates initial code $C_{0}$ based on instruction $I$. The code is executed by training on dataset $D$ to obtain results. Subsequently, AIDE iteratively derives new code $C_{1},C_{2},\dots,C_{t}$ based on the feedback.
@@
Recent research enables models to internalize the execution process, predicting test outcomes Hora (2024); team et al. (2025); Wei et al. (2025b) or assessing logic consistency directly Li et al. (2025c); Changshu et al. (2024). This capability is rigorously evaluated on reasoning-centric benchmarks Wei et al. (2025a); Gu et al. (2024a); Jain et al. (2024). Unlike traditional benchmarks that may allow for rote memorization, these tasks require models to transcend statistical pattern matching and develop a deep semantic understanding of algorithmic states and control flows Chen et al. (2025d); Sun et al. (2025a); Akhauri et al. (2025); Chen et al. (2025a), serving as the foundational capability for our proposed framework. Aligning with OpenAI’s Level 4 “Innovators” Metz (2024); Zhang et al. (2025b), this empowers agents to drive innovation by leveraging internal world models to proactively prune vast hypothesis spaces, shifting the paradigm from ensuring syntactic correctness to optimizing for semantic success. This transition resonates with the broader Data-Centric AI movement Zha et al. (2023); Cabrera et al. (2025), moving beyond model architecture to focus on the quality of evaluative signals. Specifically, our framework incorporates rationale-based preference optimization Just et al. (2024) and rigorous dataset construction criteria Shen et al. (2024) to ensure that the “implicit world model” is grounded in data-specific realities rather than abstract heuristics Tschalzev et al. (2024).
@@ Appendix B Corpus Details
To support the reproducibility of our analysis and provide a comprehensive view of the solution space, we provide detailed metadata for the corpus.
@@ B.3 Agent Evaluation Benchmark
B.4 Trajectory Sampling and Intermediate States
To clarify the composition of our Preference Corpus, it is important to note that the dataset captures the entire lifecycle of an agent’s exploration. While we filter out syntactically invalid code that crashes, our dataset is rich in logically imperfect intermediate states.
As shown in Figure 6, we generate pairs not just from the final best solution, but across the entire valid search path. This extensively tests the model’s implicit modeling capabilities to distinguish and guide improvements among messy and unfinalized intermediate code states, ensuring the agent moves from working but poor to working and good.
Root
 |-- Bug (Syntax Error) [Filtered: Handled by Interpreter]
 |    \-- Bug (Runtime Error) [Filtered]
 |         \-- Score: 0.380 (Valid, Intermediate "Half-baked") -> INCLUDED
 |              \-- Score: 0.377 (Valid, Improved) -> INCLUDED
 |                   |-- Score: 0.385 (Valid, Regressed) -> INCLUDED
 |                   |-- Score: 0.377 (Valid, Best-so-far) -> INCLUDED
 |                   \-- Score: 0.383 (Valid, Sub-optimal) -> INCLUDED
Figure 6: A representative trajectory snippet illustrating our sampling strategy. This extensively tests the model’s implicit modeling capabilities to distinguish and guide improvements among messy and unfinalized intermediate code states, ensuring the agent moves from working but poor to working and good.